12 J of work has to be done against an existing electric field to take a charge of 0.01 C from A to B. How much is the potential difference VB – VA?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Now, V base B – V base A = Potential diff = ? Charge = 0.01 C Work done = 12 J Now, Work done = Pot. Diff * Charge ⇒ Pot. Diff = 12/0.01 = 1200 Volt